OpenStudy (ny,ny):

HELP!!!!!! Solve each system using substitution. Check your solution. 3x-6y=30 y=-6x+34

OpenStudy (anonymous):

substitution basically means that you are putting something in the place of something else in the case we are going to substitute y= -6x + 34 into 3x - 6y = 30 do you understand what I am saying

OpenStudy (ny,ny):

yes i do. this is what i have so far: 3x-6(-6x+34)=30

OpenStudy (ny,ny):

but when i simplified it further, i got a really big decimal.

OpenStudy (anonymous):

so lets just look at this part first -6(-6x+34) okay

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (ny,ny):

36x - 204

OpenStudy (ny,ny):

i think.

OpenStudy (anonymous):

right now lets place it back into the rest of it 3x + 36x - 204 = 30

OpenStudy (anonymous):

so now you do 3x + 36x =

OpenStudy (ny,ny):

39x

OpenStudy (anonymous):

so now we have 39x - 204 = 30 so now add 204 to both sides

OpenStudy (ny,ny):

39x=234

OpenStudy (anonymous):

now divide by 39

OpenStudy (ny,ny):

6.

OpenStudy (anonymous):

right

OpenStudy (ny,ny):

omg, i dunno what i did wrong. thank yu so much.

OpenStudy (anonymous):

your welcome

OpenStudy (ny,ny):

ok, so that's just the x coordinate. what about y?

OpenStudy (anonymous):

oh I almost forgot

OpenStudy (ny,ny):

yeah i forgot too for a second.

OpenStudy (anonymous):

we take 6 and place it into y=-6x+34

OpenStudy (ny,ny):

ohh so we just substitute it all the way and solve for y. i think i got this. i'll ask yu the answer in a second.

OpenStudy (anonymous):

just plus 6 in x

OpenStudy (ny,ny):

yeah.

OpenStudy (ny,ny):

y=-2?

OpenStudy (anonymous):

right

OpenStudy (ny,ny):

awesome. thank yu :D

OpenStudy (anonymous):

welcome